A basketballer shoots for a goal by launching the ball at 15ms^-1 at 25° to the horizontal. What is the velocity of the ball when it is at its maximum height?
1
Expert's answer
2018-11-13T12:21:09-0500

Velocity consists of two components: v_0x and v_0y

v_0y at maximum height is zero

Therefore the velocity of the ball at its maximum height is:

v=v_0x=v_0·cos⁡α=15·cos⁡〖25 ̊〗=15·0.91=13.6 meter/sec